OpenStudy (anonymous):

Simplify: 1) 4x^2-3x-(x-x^2+3) 2) -15(4+3x)-8x=26-6(x-9). 3) 27-6(x-10)>-3x+26 write in set bulider notation

OpenStudy (anonymous):

1) 4x^2 - 3x - (x - x^2 + 3) 4x^2 - 3x - x + x^2 - 3 5x^2 - 4x - 3

OpenStudy (anonymous):

how you get 5x^2-4x-3

hartnn (hartnn):

4x^2 + x^2 =(4+1)x^2 = 5x^2

OpenStudy (anonymous):

4x^2 - 3x - x + x^2 - 3 I added 4x^2 and x^2, 5x^2 Then subtracted -3x - x -4x Then just the constant -3 hence 5x^2 - 4x - 3

OpenStudy (anonymous):

Just subsitute: is (-4)²-4(-4)-6=-6 ? 16+16-6 = -6 26 = -6 No. But if it was 4 and not (-4) it would be: 4²-4(4)-6=16-16-6=-6 So (4,-6) would be a solution.

OpenStudy (anonymous):

yeah that the solution to the question

OpenStudy (anonymous):

what question does that go to